# No Dimensional Sampling Coresets for Classification

### v1

@@ 3.2 Contraction Lemma
Before turning to the proof of this lemma, note that the only difference between this lemma and Theorem 4.12 in (Ledoux and Talagrand, 1991) is that $\varrho_{{}_{i}}$ here takes its value randomly and uniformly in $\{+\varepsilon_{i},-\varepsilon_{i}\}$ rather than $\{+1,-1\}$.
Proof [Proof of Lemma 15.] Define $\bar{\phi}_{i}(t)=\varepsilon_{i}\phi(\frac{1}{\varepsilon_{i}}t)$. Note that $\bar{\phi}_{i}(0)=0$ and $\bar{\phi}_{i}(t)$ is $1$-Lipschitz since $\phi_{i}$ is $1$-Lipschitz. Now, applying the original statement of Talagrand’s Contraction Lemma (Theorem 4.12 in (Ledoux and Talagrand, 1991)) to $\bar{\phi}_{i}$ and $\bar{\boldsymbol{T}}=\{(\varepsilon_{1}t_{1},\ldots,\varepsilon_{m}t_{m})\colon(t_{1},\ldots,t_{m})\in\boldsymbol{T}\}$ we conclude the statement as following:
$\operatorname*{\mathbb{E}}_{\varrho_{{}_{1:m}}\sim\prod_{i=1}^{m}\{\pm\varepsilon_{i}\}}\Psi\left(\frac{1}{2}\sup_{\boldsymbol{t}\in\boldsymbol{T}}\left|\sum_{i=1}^{m}\varrho_{i}\phi_{i}(t_{i})\right|\right)$ $=\operatorname*{\mathbb{E}}_{\sigma_{{}_{1:m}}\sim\prod_{i=1}^{m}\{\pm 1\}}\Psi\left(\frac{1}{2}\sup_{\bar{\boldsymbol{t}}\in\bar{\boldsymbol{T}}}\left|\sum_{i=1}^{m}\sigma_{i}\bar{\phi}_{i}(\bar{t}_{i})\right|\right)$
(Talagrand’s Contraction Lemma) $\leq\operatorname*{\mathbb{E}}_{\sigma_{{}_{1:m}}\sim\prod_{i=1}^{m}\{\pm 1\}}\Psi\left(\sup_{\bar{\boldsymbol{t}}\in\bar{\boldsymbol{T}}}\left|\sum_{i=1}^{m}\sigma_{i}\bar{t}_{i}\right|\right)$
$=\operatorname*{\mathbb{E}}_{\sigma_{{}_{1:m}}\sim\prod_{i=1}^{m}\{\pm 1\}}\Psi\left(\sup_{\boldsymbol{t}\in\boldsymbol{T}}\left|\sum_{i=1}^{m}\sigma_{i}\varepsilon_{i}t_{i}\right|\right)$
@@
$=\operatorname*{\mathbb{E}}_{\varrho_{{}_{1:m}}\sim\prod_{i=1}^{m}\{\pm\varepsilon_{i}\}}\Psi\left(\sup_{\boldsymbol{t}\in\boldsymbol{T}}\left|\sum_{i=1}^{m}\varrho_{i}t_{i}\right|\right).$
The next corollary is what we will need in our later analysis.
Corollary 16
@@
Proof  Setting $\Psi$ as identity and replacing $\phi_{i}$ with $\frac{\phi_{i}}{L}$, we conclude the result by Lemma 15.
3.3 Proof of Theorem 10
This subsection primarily focuses on proving Theorem 10 (and Theorem 5), a crucial tool enabling us to derive our main results. We start with bounding the Radamacher complexity of $\mathcal{T}_{\mathcal{F}}$.
Lemma 17
For a well-behaved measure $P$ (see Definition 8), if $s(\cdot):\mathcal{X}\longrightarrow(1,\infty)$ is an upper sensitivity function for $\mathcal{L}^{H}_{\phi,k}$ with total sensitivity $S=\int_{\mathcal{X}}s(x){\rm d}P$, then
@@ 3.3 Proof of Theorem 10
Proof  Due to similarity, we only work with $\mathcal{L}_{\phi,k}$. We first show that
$\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\left\|\sum_{i=1}^{m}\sigma_{i}\frac{\boldsymbol{K}(x_{i},\cdot)}{s(x_{i})}\right\|_{H}\leq\sqrt{\frac{m}{S}}E_{1}\quad\text{and}\quad\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\left|\sum_{i=1}^{m}\sigma_{i}\frac{1}{s(x_{i})}\right|\leq\sqrt{\frac{m}{S}}.$ (10)
@@
$\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}$ $\left\|\sum_{i=1}^{m}\sigma_{i}\frac{\boldsymbol{K}(x_{i},\cdot)}{s(x_{i})}\right\|_{H}$
$\leq\sqrt{\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\left\|\sum_{i=1}^{m}\sigma_{i}\frac{\boldsymbol{K}(x_{i},\cdot)}{s(x_{i})}\right\|^{2}_{H}}$
$=\sqrt{\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\left\langle\sum_{i=1}^{m}\sigma_{i}\frac{\boldsymbol{K}(x_{i},\cdot)}{s(x_{i})},\sum_{i=1}^{m}\sigma_{i}\frac{\boldsymbol{K}(x_{i},\cdot)}{s(x_{i})}\right\rangle_{H}}$
$=\sqrt{\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\left[\sum_{i=1}^{m}\sigma_{i}^{2}\frac{\boldsymbol{K}(x_{i},x_{i})}{s(x_{i})^{2}}+\sum_{i\neq j}\sigma_{i}\sigma_{j}\frac{\boldsymbol{K}(x_{i},x_{j})}{s(x_{i})s(x_{j})}\right]}$
$=\sqrt{\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\left[\sum_{i=1}^{m}\frac{\boldsymbol{K}(x_{i},x_{i})}{s^{2}(x_{i})}\right]}$
(since $q(x)=\frac{s(x)}{S}p(x)$) $=\frac{1}{\sqrt{S}}\sqrt{\operatorname*{\mathbb{E}}_{x_{1:m}\sim p}\left[\sum_{i=1}^{m}\frac{\boldsymbol{K}(x_{i},x_{i})}{s(x_{i})}\right]}$
(since $s(x)\geq 1$) $\leq\frac{1}{\sqrt{S}}\sqrt{\operatorname*{\mathbb{E}}_{x_{1:m}\sim p}\left[\sum_{i=1}^{m}\boldsymbol{K}(x_{i},x_{i})\right]}$
$=\frac{1}{\sqrt{S}}\sqrt{m\operatorname*{\mathbb{E}}_{x\sim p}\boldsymbol{K}(x,x)}\leq\sqrt{\frac{m}{S}}E_{1}.$
For simplicity set, $f_{w}(x)=\ell_{\phi,k}^{H}(w,x)$ and $\alpha(w)=\frac{S}{\int_{\mathcal{X}}f_{w}(x){\rm d}P(x)}$. During the proof, we need some good upper bounds for $\alpha(w)\|w\|_{H}^{2}$ and $\alpha(w)$. Note that $\int_{\mathcal{X}}f_{w}(x){\rm d}P(x)=\frac{1}{k}\|w\|_{H}^{2}+\int_{\mathcal{X}}\phi(\boldsymbol{K}(w,x)){\rm d}P(x)\geq\frac{1}{k}\|w\|_{H}^{2}$ which implies $\alpha(w)\|w\|_{H}^{2}\leq Sk$. Also,
$\int_{\mathcal{X}}f_{w}(x){\rm d}P(x)$ $=\frac{1}{k}\|w\|_{H}^{2}+\int_{\mathcal{X}}\phi(\boldsymbol{K}(w,x)){\rm d}P(x)$
@@
$\int_{\mathcal{X}}f_{w}(x){\rm d}P(x)$ $=\frac{1}{k}\|w\|_{H}^{2}+\int_{\mathcal{X}}\phi(\boldsymbol{K}(w,x)){\rm d}P(x)\geq\frac{2^{2n}}{k}$
@@
To bound the supremum within the Radamacher complexity, we utilize these inequalities to partition the parameter space. It is necessary as we want to use Lemma 16 in which we need a bounded property. Consequently,
@@
=\ = $\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{w\in\mathcal{X}}\alpha(w)\left[\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\phi(\boldsymbol{K}(w,x_{i}))+\frac{1}{k}\|w\|_{H}^{2}}{s(x_{i})}\right]$
@@
To prove (***), we deal with $M$ and $N$ separately. Note that $\bar{\phi}(\cdot)=\phi(\cdot)-\phi(0)$ is an $L$-Lipschitz functions with $\bar{\phi}(0)=0$ and thus satisfies the condition of Corollary 16. Note that
$M\leq$ $\underbrace{\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{w\in\mathcal{X}}\alpha(w)\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\boldsymbol{K}(w,x_{i}))}{s(x_{i})}\right|}_{=M_{1}}$
@@
$M_{1}^{0}=\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{\|w\|_{H}\leq 1}\alpha(w)\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\boldsymbol{K}(w,x_{i}))}{s(x_{i})}\right|$
@@
$M_{1}^{n}=\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{2^{n-1}\leq\|w\|\leq 2^{n}}\alpha(w)\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\boldsymbol{K}(w,x_{i}))}{s(x_{i})}\right|$
@@
$M_{1}^{0}$ $=\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{\|w\|_{H}\leq 1}\alpha(w)\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\boldsymbol{K}(w,x_{i}))}{s(x_{i})}\right|$
(since $\alpha(w)\leq\max(4SE_{1}k,\frac{S}{E_{2}})$) $\leq\max\left(4SE_{1}k,\frac{S}{E_{2}}\right)\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{\|w\|_{H}\leq 1}\frac{1}{m}\left|\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\boldsymbol{K}(w,x_{i}))}{s(x_{i})}\right|$
(Cor. 16 with $\varepsilon_{i}=\frac{1}{s(x_{i})}$) $\leq 2L\max\left(4SE_{1}k,\frac{S}{E_{2}}\right)\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{\|w\|_{H}\leq 1}\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\boldsymbol{K}(w,x_{i})}{s(x_{i})}\right|$
$=\frac{2L}{m}\max\left(4SE_{1}k,\frac{S}{E_{2}}\right)\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{\|w\|_{H}\leq 1}\left|\left\langle\boldsymbol{K}(w,\cdot),\sum_{i=1}^{m}\sigma_{i}\frac{\boldsymbol{K}(x_{i},\cdot)}{s(x_{i})}\right\rangle_{H}\right|$
(Cauchy–Schwarz inequality) $\leq\frac{2L}{m}\max\left(4SE_{1}k,\frac{S}{E_{2}}\right)\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{\|w\|_{H}\leq 1}\|w\|_{H}\left\|\sum_{i=1}^{m}\sigma_{i}\frac{\boldsymbol{K}(x_{i},\cdot)}{s(x_{i})}\right\|_{H}$
$\leq\frac{2L}{m}\max\left(4SE_{1}k,\frac{S}{E_{2}}\right)\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\left\|\sum_{i=1}^{m}\sigma_{i}\frac{\boldsymbol{K}(x_{i},\cdot)}{s(x_{i})}\right\|_{H}$

### v2

@@ C.2 Contraction Lemma
Before turning to the proof of this lemma, note that the only difference between this lemma and Theorem 4.12 in (Ledoux & Talagrand, 1991) is that $\varrho_{{}_{i}}$ here takes its value randomly and uniformly in $\{+\varepsilon_{i},-\varepsilon_{i}\}$ rather than $\{+1,-1\}$. However, the proof is almost identical to that of Theorem 4.12 in (Ledoux & Talagrand, 1991).
Proof of Lemma C.3.
We first prove
@@
$\operatorname*{\mathbb{E}}_{\varrho_{{}_{1:m}}\sim\prod_{i=1}^{m}\{\pm\varepsilon_{i}\}}\Psi\left(\sup_{\boldsymbol{t}\in\boldsymbol{T}}\sum_{i=1}^{m}\varrho_{{}_{i}}\phi_{i}(t_{i})\right)\leq\operatorname*{\mathbb{E}}_{\varrho_{{}_{1:m}}\sim\prod_{i=1}^{m}\{\pm\varepsilon_{i}\}}\Psi\left(\sup_{\boldsymbol{t}\in\boldsymbol{T}}\sum_{i=1}^{m}\varrho_{{}_{i}}t_{i}\right).$ (15)
To this end, it siffices to show
$\operatorname*{\mathbb{E}}_{\varrho\in\{\pm\varepsilon\}}\Psi\left(\sup_{\boldsymbol{t}\in\boldsymbol{T}}\left(t_{1}+\varrho\phi(t_{2})\right)\right)\leq\operatorname*{\mathbb{E}}_{\varrho\in\{\pm\varepsilon\}}\Psi\left(\sup_{\boldsymbol{t}\in\boldsymbol{T}}\left(t_{1}+\varrho t_{2}\right)\right),$ (16)
where $\boldsymbol{t}=(t_{1},t_{2})\subset\mathbb{R}^{2}$ is a bounded set and $\phi$ is a $1$-Lipschitz function. It suffices to prove that for all $\boldsymbol{t}=(t_{1},t_{2}),\boldsymbol{s}=(s_{1},s_{2})\in\boldsymbol{T}$, we have
$\operatorname*{\mathbb{E}}_{\varrho\in\{\pm\varepsilon\}}\Psi\left(\sup_{\boldsymbol{t}\in\boldsymbol{T}}\left(t_{1}+\varrho t_{2}\right)\right)\geq\frac{1}{2}\Psi(t_{1}+\varepsilon_{2}\phi(t_{2}))+\frac{1}{2}\Psi(s_{1}-\varepsilon_{2}\phi(s_{2}))=I.$
Since $\Psi$ is increasing, without loss of generality, we may assume that
$t_{1}+\varepsilon_{2}\phi(t_{2})\geq s_{1}+\varepsilon_{2}\phi(s_{2})\quad\quad\text{and}\quad\quad s_{1}-\varepsilon_{2}\phi(s_{2})\geq t_{1}-\varepsilon_{2}\phi(t_{2}).$ (17)
We distinguish between the following cases.
  1. 1.
    Case $\boldsymbol{t_{2},s_{2}\geq 0}$.
    • If $s_{2}\leq t_{2}$, set $a=s_{1}-\varepsilon_{2}\phi(s_{2})$, $b=s_{1}-\varepsilon_{2}s_{2}$, $a^{\prime}=t_{1}+\varepsilon_{2}t_{2}$, $b^{\prime}=t_{1}+\varepsilon_{2}\phi(t_{2})$. Since $\phi$ is 1-Lipschitz, $\phi(0)=0$, and $s_{2}\geq 0$, $|\phi(s_{2})|\leq s_{2}$ which concludes $a\geq b$. Also, by (17), it implies $b^{\prime}\geq s_{1}+\varepsilon_{2}\phi(s_{2})\geq s_{1}-\varepsilon_{2}s_{2}=b$. Furthermore, as $s_{2}\leq t_{2}$, we have $\phi(t_{2})-\phi(s_{2})\leq(t_{2}-s_{2})$ and hence
      $a-b=\varepsilon_{2}(s_{2}-\phi(s_{2}))\leq\varepsilon_{2}(t_{2}-\phi(t_{2}))=a^{\prime}-b^{\prime}.$
      Since $\Psi$ is convex and increasing, for any arbitrary fixed $x>0$, the function $\Psi(\cdot+x)-\Psi(\cdot)$ is increasing. Setting $x=a-b\geq 0$, since $b\leq b^{\prime}$, we have
      $\Psi(a)-\Psi(b)\leq\Psi(b^{\prime}+(a-b))-\Psi(b^{\prime})$
      (if $a=b$, it is true always). As $b^{\prime}+a-b\leq a^{\prime}$, we obtain $\Psi(b^{\prime}+(a-b))-\Psi(b^{\prime})\leq\Psi(a^{\prime})-\Psi(b^{\prime})$ and consequently, $\Psi(a)-\Psi(b)\leq\Psi(a^{\prime})-\Psi(b^{\prime})$ which is equivalent to $2I\leq\Psi(t_{1}+\varepsilon_{2}t_{2})+\Psi(s_{1}-\varepsilon_{2}s_{2}).$
    • If $t_{2}\leq s_{2}$, set $a=t_{1}+\varepsilon_{2}\phi(t_{2})$, $b=t_{1}-\varepsilon_{2}t_{2}$, $a^{\prime}=s_{1}+\varepsilon_{2}s_{2}$, $b^{\prime}=s_{1}-\varepsilon_{2}\phi(s_{2})$. Since $\phi$ is $1$-Lipschitz, $|\phi(t_{2})|\leq t_{2}$ which implies $a\geq b$. Again, by (17),
      $b^{\prime}=s_{1}-\varepsilon_{2}\phi(s_{2})\geq t_{1}-\varepsilon_{2}\phi(t_{2})\geq t_{1}-\varepsilon_{2}t_{2}=b.$
      As $\phi(t_{2})-\phi(s_{2})\leq s_{2}-t_{2}$,
      $a-b=\varepsilon_{2}(t_{2}+\phi(t_{2}))\leq\varepsilon_{2}(s_{2}+\phi(s_{2}))=a^{\prime}-b^{\prime}.$
      Again, as $\Psi(\cdot+x)-\Psi(\cdot)$ is increasing for each fixed $x>0$ and $b\leq b^{\prime}$, by setting $x=a-b\geq 0$, we have
      $\Psi(a)-\Psi(b)\leq\Psi(b^{\prime}+(a-b))-\Psi(b^{\prime})$
      (if $a=b$, it is true always). As $b^{\prime}+a-b\leq a^{\prime}$, we obtain $\Psi(b^{\prime}+(a-b))-\Psi(b^{\prime})\leq\Psi(a^{\prime})-\Psi(b^{\prime})$ and consequently, $\Psi(a)-\Psi(b)\leq\Psi(a^{\prime})-\Psi(b^{\prime})$ which is equivalent to $2I\leq\Psi(s_{1}+\varepsilon_{2}s_{2})+\Psi(t_{1}-\varepsilon_{2}t_{2}).$
  2. 2.
    Case $\boldsymbol{t_{2},s_{2}\leq 0}$. This case is similar to the previous case.
  3. 3.
    Case $\boldsymbol{t_{2}\geq 0,s_{2}\leq 0}$. Since $\Psi$ is increasing and $\phi(t_{2})\leq t_{2},-\phi(s_{2})\leq-s_{2}$, we have
    $2I\leq\Psi(t_{1}+\varepsilon_{2}t_{2})+\Psi(s_{1}-\varepsilon_{2}s_{2}).$
  4. 4.
    Case $\boldsymbol{t_{2}\leq 0,s_{2}\geq 0}$. This case follows with a similar argument to Case (3). This completes the proof of (16) and consequently the proof of (15).
To conclude the theorem, by convexity of $\Psi$,
$\operatorname*{\mathbb{E}}_{\varrho_{{}_{1:m}}\sim\prod_{i=1}^{m}\{\pm\varepsilon_{i}\}}\Psi\left(\frac{1}{2}\sup_{\boldsymbol{t}\in\boldsymbol{T}}\left|\sum_{i=1}^{m}\varrho_{{}_{i}}\phi_{i}(t_{i})\right|\right)\leq$ $~{}\frac{1}{2}\operatorname*{\mathbb{E}}_{\varrho_{{}_{1:m}}\sim\prod_{i=1}^{m}\{\pm\varepsilon_{i}\}}\Psi\left(\sup_{\boldsymbol{t}\in\boldsymbol{T}}\left[\sum_{i=1}^{m}\varrho_{{}_{i}}\phi_{i}(t_{i})\right]^{+}\right)$
$+\frac{1}{2}\operatorname*{\mathbb{E}}_{\varrho_{{}_{1:m}}\sim\prod_{i=1}^{m}\{\pm\varepsilon_{i}\}}\Psi\left(\sup_{\boldsymbol{t}\in\boldsymbol{T}}\left[\sum_{i=1}^{m}\varrho_{{}_{i}}\phi_{i}(t_{i})\right]^{-}\right)$
[…]
=\ = $\operatorname*{\mathbb{E}}_{\varrho_{{}_{1:m}}\sim\prod_{i=1}^{m}\{\pm\varepsilon_{i}\}}\Psi\left(\left[\sup_{\boldsymbol{t}\in\boldsymbol{T}}\sum_{i=1}^{m}\varrho_{{}_{i}}\phi_{i}(t_{i})\right]^{+}\right)$
[…]
where (*) is true by applying (15) on the convex and increasing function $\Psi(\max(0,\cdot))$. This completes the proof. ∎
Corollary C.4.
@@
Proof.
Setting $\Psi$ as identity and replacing $\phi_{i}$ with $\frac{\phi_{i}}{L}$, we conclude the result by Lemma C.3. ∎
In our setting, we only deal with the case that each $\varepsilon_{i}\in(0,1]$. This particular case can be directly inferred from Theorem 4.12 in (Ledoux & Talagrand, 1991). Nevertheless, Lemma C.3 and Corollary C.4, in the presented general form, can be of independent interest.
C.3 Proof of Theorem A.3
This subsection primarily focuses on proving Theorem A.3 (and Theorem 1.4), a crucial tool enabling us to derive our main results. We start with bounding the Radamacher complexity of $\mathcal{T}_{\mathcal{F}}$.
Lemma C.5.
For a well-behaved measure $P$ (see Definition A.1), if $s(\cdot):\mathcal{X}\longrightarrow(1,\infty)$ is an upper sensitivity function for $\mathcal{L}^{H}_{\phi,k}$ with total sensitivity $S=\int_{\mathcal{X}}s(x){\rm d}P$, then
@@
Proof.
Due to similarity, we only work with $\mathcal{L}_{\phi,k}$. We first show that
$\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\left\|\sum_{i=1}^{m}\sigma_{i}\frac{\mathsf{K}(x_{i},\cdot)}{s(x_{i})}\right\|_{H}\leq\sqrt{\frac{m}{S}}E_{1}\quad\text{and}\quad\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\left|\sum_{i=1}^{m}\sigma_{i}\frac{1}{s(x_{i})}\right|\leq\sqrt{\frac{m}{S}}.$ (18)
@@
$\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}$ $\left\|\sum_{i=1}^{m}\sigma_{i}\frac{\mathsf{K}(x_{i},\cdot)}{s(x_{i})}\right\|_{H}$
$\leq\sqrt{\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\left\|\sum_{i=1}^{m}\sigma_{i}\frac{\mathsf{K}(x_{i},\cdot)}{s(x_{i})}\right\|^{2}_{H}}$
$=\sqrt{\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\left\langle\sum_{i=1}^{m}\sigma_{i}\frac{\mathsf{K}(x_{i},\cdot)}{s(x_{i})},\sum_{i=1}^{m}\sigma_{i}\frac{\mathsf{K}(x_{i},\cdot)}{s(x_{i})}\right\rangle_{H}}$
$=\sqrt{\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\left[\sum_{i=1}^{m}\sigma_{i}^{2}\frac{\mathsf{K}(x_{i},x_{i})}{s(x_{i})^{2}}+\sum_{i\neq j}\sigma_{i}\sigma_{j}\frac{\mathsf{K}(x_{i},x_{j})}{s(x_{i})s(x_{j})}\right]}$
$=\sqrt{\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\left[\sum_{i=1}^{m}\frac{\mathsf{K}(x_{i},x_{i})}{s^{2}(x_{i})}\right]}$
(since $q(x)=\frac{s(x)}{S}p(x)$) $=\frac{1}{\sqrt{S}}\sqrt{\operatorname*{\mathbb{E}}_{x_{1:m}\sim p}\left[\sum_{i=1}^{m}\frac{\mathsf{K}(x_{i},x_{i})}{s(x_{i})}\right]}$
(since $s(x)\geq 1$) $\leq\frac{1}{\sqrt{S}}\sqrt{\operatorname*{\mathbb{E}}_{x_{1:m}\sim p}\left[\sum_{i=1}^{m}\mathsf{K}(x_{i},x_{i})\right]}$
$=\frac{1}{\sqrt{S}}\sqrt{m\operatorname*{\mathbb{E}}_{x\sim p}\mathsf{K}(x,x)}\leq\sqrt{\frac{m}{S}}E_{1}.$
For simplicity set, $f_{w}(x)=\ell_{\phi,k}^{H}(w,x)$ and $\alpha(w)=\frac{S}{\int_{\mathcal{X}}f_{w}(x){\rm d}P(x)}$. During the proof, we need some good upper bounds for $\alpha(w)\|w\|_{H}^{2}$ and $\alpha(w)$. Note that $\int_{\mathcal{X}}f_{w}(x){\rm d}P(x)=\frac{1}{k}\|w\|_{H}^{2}+\int_{\mathcal{X}}\phi(\mathsf{K}(w,x)){\rm d}P(x)\geq\frac{1}{k}\|w\|_{H}^{2}$ which implies $\alpha(w)\|w\|_{H}^{2}\leq Sk$. Also,
$\int_{\mathcal{X}}f_{w}(x){\rm d}P(x)$ $=\frac{1}{k}\|w\|_{H}^{2}+\int_{\mathcal{X}}\phi(\mathsf{K}(w,x)){\rm d}P(x)$
@@
$\int_{\mathcal{X}}f_{w}(x){\rm d}P(x)$ $=\frac{1}{k}\|w\|_{H}^{2}+\int_{\mathcal{X}}\phi(\mathsf{K}(w,x)){\rm d}P(x)\geq\frac{2^{2n}}{k}$
@@
To bound the supremum within the Radamacher complexity, we utilize these inequalities to partition the parameter space. It is necessary as we want to use Lemma C.4 in which we need a bounded property. Consequently,
@@
=\ = $\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{w\in\mathcal{X}}\alpha(w)\left[\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\phi(\mathsf{K}(w,x_{i}))+\frac{1}{k}\|w\|_{H}^{2}}{s(x_{i})}\right]$
@@
To prove (*), we deal with $M$ and $N$ separately. Note that $\bar{\phi}(\cdot)=\phi(\cdot)-\phi(0)$ is an $L$-Lipschitz functions with $\bar{\phi}(0)=0$ and thus satisfies the condition of Corollary C.4. Note that
$M\leq$ $\underbrace{\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{w\in\mathcal{X}}\alpha(w)\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\mathsf{K}(w,x_{i}))}{s(x_{i})}\right|}_{=M_{1}}$
@@
$M_{1}^{0}=\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{\|w\|_{H}\leq 1}\alpha(w)\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\mathsf{K}(w,x_{i}))}{s(x_{i})}\right|$
@@
$M_{1}^{n}=\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{2^{n-1}\leq\|w\|\leq 2^{n}}\alpha(w)\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\mathsf{K}(w,x_{i}))}{s(x_{i})}\right|$
@@
$M_{1}^{0}$ $=\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{\|w\|_{H}\leq 1}\alpha(w)\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\mathsf{K}(w,x_{i}))}{s(x_{i})}\right|$
(since $\alpha(w)\leq\max(4SE_{1}k,\frac{S}{E_{2}})$) $\leq\max\left(4SE_{1}k,\frac{S}{E_{2}}\right)\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{\|w\|_{H}\leq 1}\frac{1}{m}\left|\sum_{i=1}^{m}\sigma_{i}\frac{\bar{\phi}(\mathsf{K}(w,x_{i}))}{s(x_{i})}\right|$
(Cor. C.4 with $\varepsilon_{i}=\frac{1}{s(x_{i})}$) $\leq 2L\max\left(4SE_{1}k,\frac{S}{E_{2}}\right)\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{\|w\|_{H}\leq 1}\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\frac{\mathsf{K}(w,x_{i})}{s(x_{i})}\right|$
$=\frac{2L}{m}\max\left(4SE_{1}k,\frac{S}{E_{2}}\right)\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{\|w\|_{H}\leq 1}\left|\left\langle\mathsf{K}(w,\cdot),\sum_{i=1}^{m}\sigma_{i}\frac{\mathsf{K}(x_{i},\cdot)}{s(x_{i})}\right\rangle_{H}\right|$
(Cauchy–Schwarz inequality) $\leq\frac{2L}{m}\max\left(4SE_{1}k,\frac{S}{E_{2}}\right)\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\sup_{\|w\|_{H}\leq 1}\|w\|_{H}\left\|\sum_{i=1}^{m}\sigma_{i}\frac{\mathsf{K}(x_{i},\cdot)}{s(x_{i})}\right\|_{H}$
$\leq\frac{2L}{m}\max\left(4SE_{1}k,\frac{S}{E_{2}}\right)\operatorname*{\mathbb{E}}_{x_{1:m}\sim q}\operatorname*{\mathbb{E}}_{\sigma_{1:m}\sim\{-1,1\}}\left\|\sum_{i=1}^{m}\sigma_{i}\frac{\mathsf{K}(x_{i},\cdot)}{s(x_{i})}\right\|_{H}$